Circular No. 6232

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1995 S1 = P/1846 D1 (DE VICO) Preliminary orbital elements, from 39 observations Sept. 17–20, bear

a striking resemblance to those of D/1846 D1 (1846 IV). Though the predictions were known to have an uncertainty of several years, R. J. Buckley (1979, JBAA 89(3), 260) had calculated T = 1996 July 3. It has not been possible to link the two apparitions, despite great effort to do so, probably because of nongravitational forces that require a longer arc of observation at the current return. The following orbital elements by the undersigned were computed from an unpublished 1846 orbit utilizing 17 observations 1846 Feb. 27–Apr. 27 (mean residual 2′′.9); the observations had been rereduced recently by B. G. Marsden using the new PPM Star Catalogue and a computer program written by G. V. Williams. The resulting orbit was adjusted (with full perturbations) so that e in 1846 produced an orbit with T ∼ 1995 Oct. 6, which is the value found by performing a 2-body calculation using the 1995 observations alone; the 1995-epoch elements from the 1846 observations were then differentially corrected (2-body calculation) while holding e and q fixed (and solving for T and the angles):

T = 1995 Oct. 6.026 TT  ω = 12.^{o}978
e = 0.96274  Ω = 79.626 }2000.0
q = 0.65891 AU  i = 85.382
a = 17.68245 AU n^{o} = 0.013255 P = 74.36 years

These elements suggest T ∼ 1922 Apr. for the comet’s last (missed) return to perihelion. Crommelin (1917, JBAA 27, 265; 1920, JBAA 30,

135) suggested that the most probable T at its first predicted return was around 1921 Nov., though search ephemerides were provided for several years. Ephemeris from the above elements:

1995TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Sept.20 9^{h}11.^{m}46 − 0^{o}19.^{′}5 1.157 0.737 39.^{o}1 59.^{o}2 6.5
22 9 19.30 + 1 42.4 1.129 0.720 38.9 61.1 6.3
24 9 27.80 + 3 50.1 1.102 0.704 38.7 63.0 6.2
26 9 37.07 + 6 03.3 1.076 0.691 38.6 64.9 6.1
28 9 47.19 + 8 21.5 1.052 0.680 38.5 66.8 5.9
30 9 58.24 +10 43.4 1.030 0.671 38.5 68.5 5.8

1995 September 20 Daniel W. E. Green